OpenStudy (jhannybean):

Determine whether the value of the gas constant R would be: a) higher, b) lower, c) cause no change if: Zn(s) contains contaminants that are insoluble with HCl solution (hint: follow a setup for the calculation of R to see the effect of each error)

OpenStudy (jhannybean):

@Photon336

OpenStudy (jhannybean):

So I started with the reaction: \(\sf Zn~(s) ~+~ HCl ~(aq)~\rightarrow ~ZnCl_2~(aq)~+~H_2~(g)~\) the ideal gas law only applies to gasses, therefore setting up my ratio for R, I get, \(\sf R=\dfrac{P_{H_2}V_{H_2}}{n_{H_2}T_{H_2}}\)

OpenStudy (jhannybean):

Does the contamination of zinc \(\mathbf {lower}\) the actual amount of zinc reacting with the HCl?

OpenStudy (jhannybean):

\[\sf Zn~(s) ~+~ \color{red}{2}HCl ~(aq)~\rightarrow ~ZnCl_2~(aq)~+~H_2~(g)~\]

OpenStudy (anonymous):

Contamination would lower the amount of zinc before the reaction or at the reaction.

OpenStudy (jhannybean):

Ohhh, alright!

OpenStudy (anonymous):

And they're insoluble with HCl, which leaves a lighter zinc.

OpenStudy (jhannybean):

Then how would using a lower amount of zinc effect the volume of \(\sf H_2\) ?

OpenStudy (photon336):

I agree with @Shalante

OpenStudy (anonymous):

Not 100% sure though.

OpenStudy (anonymous):

I would say Zinc+contaminants+HCl>> Zinc Chloride+Hydrogen gas+contaminants. Mass of Contaminants before=mass of contaminants after.

OpenStudy (jhannybean):

Oh... the way I calculated it I think I might have done something wrong. I took the mass of zinc I had measured out, and converted that into moles, used PV=nRT to find the volume with the respective moles f zinc from the measured amount. Then I compared it to zinc with contaminants, but the mass of zinc with contaminants should have been lower, and not a higher amount I think.

OpenStudy (jhannybean):

So.. i'll write out what I did.

OpenStudy (anonymous):

PV=nRT on a solid? (zinc)

OpenStudy (jhannybean):

\[\text{mass of zinc w/o contamin.} =0.1901~g~Zn\]\[\sf ~P_{H_2} =~0.961~atm~\\ V_{H_2}~=~?~\\ n_{H_2} = 0.1901~g~Zn ~\times~\frac{1~mol~Zn}{65.39~g~Zn} = 0.002907~mol~Zn \]\[\sf PV=nRT \\ (0.961~ atm)V_{H_2} =~ (0.002907~ mol)\left(\frac{0.0821~L ~atm}{~ mol~ K}\right) (295~K) \\ V_{H_2} = 0.0732~ L\]

OpenStudy (jhannybean):

Ack. nvm. I had to change to moles of \(H_2\)

OpenStudy (jhannybean):

\[\sf ~P_{H_2} =~0.961~atm~\\ V_{H_2}~=~?~\\ n_{H_2} = 0.1901~g~Zn ~\times~\frac{1~mol~Zn}{65.39~g~Zn}~\times~\frac{1~mol~H_2}{1~mol~Zn } = 0.002097 ~mol~H_2 ~\]

OpenStudy (jhannybean):

I did the same thing by increasing the g of zinc with contaminants, but i think that is wrong to do because it will give me a higher value volume of H2.

OpenStudy (anonymous):

Grams of zinc decreases with contaminants . That is my opinion.

OpenStudy (anonymous):

But zinc can be a contaminant lol. If zinc contains contaminants, it cant have more zinc. Either it stays the same or decreases. Thats what I think.

OpenStudy (jhannybean):

yeah, the correct answer was that it \(\mathbf {lowered}\) the value of the gas constant, which would only make sense if the \(\mathbf{volume}\) of the \(\sf H_2\) gas was \(\mathbf{decreased}\) due to the \(\mathbf{decrease}\) in moles of \(\sf H_2\) produced by a \(\mathbf{decrease}\) in the moles of contaminated Zn.

OpenStudy (anonymous):

Oh yeah, I mentioned that earlier and photon agree. Because I think contaminant is in some case can decrease an element by removing it through chemical reaction. Just like how soap kills bacteria, but contamination is the opposite.

OpenStudy (anonymous):

I only studied contamination in groundwater and the amount of contamination that adds up to it makes the water undrinkable due to lack to hydrogen and oxygen.

OpenStudy (anonymous):

I didnt realize my drawing was cut out. I used conservation of mass on a table earlier, but it was cut off. I didnt realize. It showed that the grams of hydrogen gas decrease in mass with contaminants

OpenStudy (jhannybean):

Ahh I see I see. It makes sense now!

OpenStudy (jhannybean):

Thank you @Shalante :)

OpenStudy (jhannybean):

I might have some other questions later!

OpenStudy (anonymous):

Ok, I'll check it out later. See you soon.

OpenStudy (jhannybean):

^^